A farmer has 400 acres of available land and $80,000 to spend He wants to plant the combination of crops which maximizes hisPlease show work using matrices, thank you!
Answer #1

Dual of the LP

Minimize W = 400y1+80000y2

S.T.

y1+400y2>=180

y1+160y2>=60

y1+280y2>=90

y1,y2>=0

b) Sensitivity Report and shadow prices

. 2

The sensitivity report says the shadow price of constraint 1 is 0 and constraint 2 is 0.45

Making the change in land acres will not affect the optimal value but making one unit change in budget will change the optimal value by 0.45

So changing acres to 360 and budget to 84000

Change in profit = 0+0.45*4000=1800

New profit = 36000+1800=$37800

c) New optimal plan

The acre of land of potatoes = 190

Acre of land of corn = 0

Acre of land of cabbage =0

Total profit = 34200